OpenStudy (anonymous):

1. Identify the sampling method. You want to find how many students use public transportation. You interview every fifth teenager you see exiting a movie theater. A. random B. systematic << C. stratified D. none of these right?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

systematic is a basic pattern or rule

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

The two box-and-whisker plots below show the times in seconds for two teams in a 100 m dash. What do the interquartile ranges tell you about the two teams? A. Team A has more consistent times B. Team B has more consistent times << C. Overall team A is faster than team B D. Overall team B is faster than team A https://study.ashworthcollege.edu/access/content/group/38441f94-71d6-46e6-8dd7-ece73f5a6548/algebra1_exam_12_files/mc009-1.jpg

jimthompson5910 (jim_thompson5910):

team B is not as spread out as team A, so that would suggest there is not as much deviation (and it would mean team B is more consistent) so I agree

OpenStudy (anonymous):

3. The back-to-back stem-and-leaf plot below show exam scores from two different math classes. Which class has a greater mean score? Which class has a greater median score? Class A Class B 1 5 4 2 1 6 8 5 4 5 7 7 9 6 1 6 6 6 7 8 9 7 2 5 6 6 1 2 8 0 0 4 8 9 1 9 3 5 6 7 A. greater mean = class A greater median = class A B. greater mean = class B greater median = class B C. greater mean = class B greater median = class A D. greater mean = class A greater median = class B is it B? cuz i got C wrong?

jimthompson5910 (jim_thompson5910):

what did you get for the mean of class A

OpenStudy (anonymous):

class a = 1, 5, 1, 6, 8, 5, 7, 7, 9,6,6,7,8,9,1,2,1 class b = 2,4,1,6,2,5,6,60,0,4,8,9,3,5,6,7 the ones in the middle =4,5,6,7,8,9

OpenStudy (anonymous):

class a=5.23

jimthompson5910 (jim_thompson5910):

I'm getting 68.176 for the mean of class A

jimthompson5910 (jim_thompson5910):

I'm getting these values for class A 41,45,51,56,58,65,67,67,69,76,76,77,78,79,81,82,91

jimthompson5910 (jim_thompson5910):

you average this list: 41,45,51,56,58,65,67,67,69,76,76,77,78,79,81,82,91 NOT this list 1, 5, 1, 6, 8, 5, 7, 7, 9,6,6,7,8,9,1,2,1 to get the mean for class A

jimthompson5910 (jim_thompson5910):

Mean = (Sum of the values)/(Number of values) Mean = (41+45+51+56+58+65+67+67+69+76+76+77+78+79+81+82+91)/17 Mean = 1159/17 Mean = 68.176

OpenStudy (anonymous):

ooooh, now i get it...

jimthompson5910 (jim_thompson5910):

do the same for class B and compare the two means

jimthompson5910 (jim_thompson5910):

after you find the means, you'll need to compare the medians as well

OpenStudy (anonymous):

so, class b= 42,54,61,66,72,75,76,76,80,80,84,88,89,93,95,96,97/17 mean= 77.88

jimthompson5910 (jim_thompson5910):

good, you got the correct mean

jimthompson5910 (jim_thompson5910):

so class B has the greater mean

jimthompson5910 (jim_thompson5910):

now compare the medians

jimthompson5910 (jim_thompson5910):

oh I guess since you know it's not C, then it has to be B since that's the only one left where class B has the greater mean

OpenStudy (anonymous):

class a =69 class b= 80?

jimthompson5910 (jim_thompson5910):

one sec and let me confirm

OpenStudy (anonymous):

sure!

jimthompson5910 (jim_thompson5910):

good on both, so we can see that class B also has the greater median

OpenStudy (anonymous):

Thank youuu for the helpp!!! alright, last question. sorry. The two box-and-whisker plots below show the scores on a math exam for two classes. What do the interquartile ranges tell you about the two classes? A. Class A has more consistent scores B. Class B has more consistent scores C. Overall class A performed better than class B D. Overall class B performed better than class A https://study.ashworthcollege.edu/access/content/group/38441f94-71d6-46e6-8dd7-ece73f5a6548/algebra1_exam_12_files/mc010-1.jpg https://study.ashworthcollege.edu/access/content/group/38441f94-71d6-46e6-8dd7-ece73f5a6548/algebra1_exam_12_files/mc010-2.jpg

OpenStudy (anonymous):

i say C am i right?/ :)

jimthompson5910 (jim_thompson5910):

hmm not 100% sure, but I'd say either C (since A does have a lot of scores higher than B) or B (since B has a smaller IQR)

OpenStudy (anonymous):

Thank youu!

jimthompson5910 (jim_thompson5910):

you're welcome